OpenStudy (babyslapmafro):

The anti-derivative of tan(x) is -log(cosx). How is that determined? Is it a mathematical fact that I must memorize or is there a method to obtain this if you are given the integral of tan(x)dx?

OpenStudy (abb0t):

\[\tan = \frac{ \sin }{ \cos }\]

OpenStudy (babyslapmafro):

oh right...thanks